[Brahms-l] Shift report 20040304 16:00-24:00

From: Apache <apache@bnl.gov>
Date: Thu Mar 04 2004 - 00:18:11 EST
------------------------------------------------------------
Date: Thu Mar  4 00:18:11 2004, Shift: 16:00-24:00
Supervisor: pj, with: 
Log book#15 pages 72 - 72

Shift Summary:
------------------------------------------------------------
1. Access problems plagued the shift.  But in the end we 
got some beam.

2. Nothing. But run 10742 still running with: FS 8deg
full B, MRS 90deg 1/2 B

3. TPM1 tripped as I turned on HV.  After turning it off
and on, it worked like a charm.
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Thu Mar 4 00:18:23 2004

This archive was generated by hypermail 2.1.8 : Thu Mar 04 2004 - 00:18:43 EST